OpenStudy (anonymous):

Solve this system of equations: 2x - 5y =1 4x + 2y = 14

hero (hero):

2x-5y = 1 4x+2y = 14 y = (1-2x)/-5 y = (14-4x)/2 y = y 2(1-2x) = -5(14-4x) 2 -4x= -70+20x 72 = 24x 72/24 = x x = 3 y = 1

OpenStudy (anonymous):

Whay do you multiply it my -5 and 2?

hero (hero):

Because I crossed multiplied by fractions I created once I isolated y

OpenStudy (anonymous):

why do you cross multiply?

hero (hero):

Because I set two fractions equal to one another. I set them equal to each other because they are expressions that represent intersecting lines. The goal is to find the (x,y) point where they intersect. Setting the expressions equal to each other is and finding the (x,y) points is like graphing both lines to find the point, except one is an algebraic method and the other is a graphic method.

OpenStudy (anonymous):

oh, thank you. :D